# Fundamental multiplication formulas

### v1

@@ 2. The ${\boldsymbol{q}}$-Schur algebras of type $D$: an algebraic setting
$S_{{\boldsymbol{q}}}^{\kappa}(n,r)=\text{\rm End}_{\check{H}}\big{(}\bigoplus_{\lambda\in\Lambda(n+1,r)}x_{\lambda}H_{{\boldsymbol{q}},1}|_{\check{H}}\big{)}$ (2.1.1)
Let ${\mathcal{Z}}:=\mathbb{Z}[\upsilon,\upsilon^{-1}]$. By base change to ${\mathcal{Z}}$, we obtain the Hecke algebra ${\mathcal{H}},\check{\mathcal{H}}$ and the $\upsilon$-Schur algebras
${\mathcal{H}}:=H_{{\boldsymbol{q}},1}\otimes{\mathcal{Z}},\;\;\check{\mathcal{H}}:=\check{H}\otimes{\mathcal{Z}},\;\;{\mathcal{S}}_{\upsilon,1}^{\textsc{b}}(n,r):=S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)\otimes_{{\mathcal{A}}}{\mathcal{Z}},\;\;{\mathcal{S}}_{\upsilon}^{\kappa}(n,r)=S_{{\boldsymbol{q}}}^{\kappa}(n,r)\otimes_{{\mathcal{A}}}{\mathcal{Z}}.$ (2.1.2)
@@
Let ${\mathfrak{S}}_{\mu}$ be the Young subgroup of ${\mathfrak{S}}_{2r}$ associated with $\mu$. This is the stabiliser subgroup of the standard Young tabloid [^2] If we identify $\mu$ with its Young diagram consisting of $\mu_{1}$ boxes in row 1, and $\mu_{2}$ boxes in row 2 and so on, then a $\mu$-tableau is obtained by filling the numbers $1,2,\ldots,2r$ into boxed. A $\mu$-tabloid is a row-equivalent class of $\mu$-tableaux. $R^{\mu}=R^{\mu}_{1}\times\cdots\times R_{n}^{\mu}$, where $R^{\mu}_{i}:=[\widetilde{\mu}_{i-1}+1,\widetilde{\mu}_{i-1}+\mu_{i}]$, for all $i=1,2,\ldots,N$. In other words,
@@
consisting of permutations which stabilise each of the subsets $R^{\mu}_{i}$.
@@
^$\widehat{\ }:\Lambda(n+1,r)\longrightarrow\Lambda(2n+1,2r),\;\;\lambda\longmapsto\widehat{\lambda},$ (2.1.8)
and denote its image by $\widehat{\Lambda}(n+1,r)$. We now describe the parabolic subgroup $W_{\lambda}$ defined in (2.0.7) in terms of stabilisers of the subsets $R_{i}^{\lambda}:=R_{i}^{\widehat{\lambda}}$, for $1\leq i\leq{\mathsf{m}}$, where ${\mathsf{m}}={\mathsf{m}}(\lambda)$ denotes the maximal index $i$ satisfying $\lambda_{i}\neq 0$.
@@
    Suppose the double coset ${\mathfrak{S}}_{\widehat{\lambda}}w{\mathfrak{S}}_{\widehat{\mu}}$ defines a $(2n+1)\times(2n+1)$ matrix $A=(a_{i,j})$ over $\mathbb{N}$ whose entries sum to $2r$. Then ${\mathfrak{S}}_{\widehat{\lambda}}w{\mathfrak{S}}_{\widehat{\mu}}$ is stabilised by $\theta$ if and only if $a_{i,j}=a_{N+1-i,N+1-j}$ for all $i,j\in[1,N]$ ($N=2n+1$). In particular, for ${\widehat{d}}=\iota(d)\in{\mathfrak{S}}_{2r}$, $\theta$ stabilises ${\mathfrak{S}}_{\widehat{\lambda}}{\widehat{d}}{\mathfrak{S}}_{\widehat{\mu}}$ and the double coset $W_{\lambda}dW_{\mu}=({\mathfrak{S}}_{\widehat{\lambda}}{\widehat{d}}{\mathfrak{S}}_{\widehat{\mu}})^{\theta}$.
@@
$\text{\rm ro}(A)=\Big{(}\sum_{j=1}^{N}a_{1,j},\sum_{j=1}^{N}a_{2,j},\ldots,\sum_{j=1}^{N}a_{N,j}\Big{)},\quad\text{\rm ro}(A)=\Big{(}\sum_{i=1}^{N}a_{i,1},\sum_{i=1}^{N}a_{i,2},\ldots,\sum_{i=1}^{N}a_{i,N}\Big{)}.$ (2.3.2)
@@
Each triple $(\lambda,d,\mu)\in{\mathcal{D}}(n,r)$ defines a standard basis element in $S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)$. We denote this element by $e_{A}$ if $A=\mathfrak{d}(\lambda,d,\mu)$. By definition,
@@
The set $\{e_{A}\mid A\in\Xi_{2n+1,2r}\}$ forms a basis for $S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)$ which we call the standard basis.
We now extend the lemma above to parabolic subgroups of ${\check{W}}$; compare [19]. Let $\lambda\in\Lambda(n+1,r)$ have maximal index ${\mathsf{m}}$.
@@
(1) The case (2) above shows that distinct tabloids may have the same stabiliser subgroups. This phenomenon is similar to the type $A$ fact that distinct compositions may define the same Young subgroup of the symmetric group.
(2) Consider the parabolic subgroup $W_{\mu}=\langle s_{1},\ldots,s_{r-1}\rangle$ of $W$, where $\mu={(r,0^{n})}$. This is also a parabolic subgroup of $\check{W}$. That is $W_{\mu}=\check{W}_{\mu}$. Clearly, $\varsigma=s_{r}s_{r-1}s_{r}\in{\mathcal{D}}_{\mu,\mu}$, but $s_{r-1}=s_{r}ds_{r}$ is not distinguished in the double coset $\check{W}_{\mu}(s_{r}ds_{r})\check{W}_{\mu}=\check{W}_{\mu}$.
We are now ready to introduce the $q$-Schur algebra of type $D$ (see [19, (A.3.3)] and compare [11, Def. 2.5]). Let
@@
$S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$ $:=\text{\rm End}_{\check{H}}\Big{(}\bigoplus_{\alpha\in\Lambda^{\textsc{d}}(n,r)}{\check{x}}_{\alpha}{\check{H}}\Big{)}\quad(\text{over }{\mathcal{A}}=\mathbb{Z}[{\boldsymbol{q}}]);$
${\mathcal{S}}_{\upsilon}^{\textsc{d}}(n,r)$ $:=\text{\rm End}_{\check{\mathcal{H}}}\Big{(}\bigoplus_{\alpha\in\Lambda^{\textsc{d}}(n,r)}{\check{x}}_{\alpha}{\check{\mathcal{H}}}\Big{)}\quad(\text{over }{\mathcal{Z}}=\mathbb{Z}[\upsilon,\upsilon^{-1}]).$
@@
In [4], a geometric setting is given to the $q$-Schur algebra $S^{\jmath}_{q}(n,r)$ of type $B$ in equal parameter. We now modify their construction using $(2n+1)$-step filtrations for the $2r$-dimensional space $\mathbb{F}_{q}^{2r}$ (see [14]) and the action by the orthogonal group ${\text{\rm O}}_{2r}(q)$. This results in a $q$-Schur algebra in unequal parameters. Restricting the action to the subgroup ${\text{\rm SO}}_{2r}(q)$ will lead to a geometric setting for the $q$-Schur algebra of type $D$; see Theorems 3.10 and 4.15 in the next two sections.
@@ 3.1. A geometric setting for the endomorphism algebra of a permutation module
The $G$-action extends diagonally to $X\times X$. Consider three projection maps $\pi_{i,j}:X\times X\times X\to X\times X$ sending $(x_{1},x_{2},x_{3})$ to $(x_{i},x_{j})$ for $(i,j)\in\{(1,2),(2,3),(1,3)\}$. Then $\mathfrak{F}_{G}(X\times X,{\mathcal{R}})$ becomes an associative algebra with multiplication defined by $f*f^{\prime}=(\pi_{1,3})_{!}(\pi_{1,2}^{*}f\cdot\pi_{2,3}^{*}f^{\prime})$. In other words,
@@ 3.2. Finite orthogonal groups and flag varieties
We now look at the case, where $G=O_{2r}(q)$ or $SO_{2r}$ is the finite orthogonal or special orthogonal group, and $X=\mathcal{F}_{n,r}^{\jmath}$ is the set of $n$-step filtrations (or partial flag variety) of isotropic subspaces.
@@
This is a group with a split BN-pair, where $B=\widehat{B}(q)\cap{\text{\rm O}}_{2r}(q)$ with $\widehat{B}(q)$ being the Borel subgroup of ${\text{\rm GL}}_{2r}(q)$ consisting upper triangular matrices and $N=\{\dot{w}\mid w\in W\}$ (see, e.g., [15, p.80]), where
@@ 3.3. Parabolic subgroups of ${\text{\rm O}}_{2r}(q)$
Consider the standard basis $e_{1},e_{2},\ldots,e_{2r}$ for $\mathbb{F}_{q}^{2r}$ and form the (standard) pure isotropic complete flag $\mathsf{F}^{s}$ with
@@
Then the subgroup $B=\text{Stab}_{G(q)}(\mathsf{F})$ is a Borel subgroup of $G=G(q):={\text{\rm O}}_{2r}(q)$. Note that if $\widehat{B}(q)$ denotes the Borel subgroup of ${\text{\rm GL}}_{2r}(q)$ consisting upper triangular matrices, then $B(q)=\widehat{B}(q)\cap G(q)$. Moreover, the mapping $G(q)\to\mathcal{B}_{r}$ sending $g$ to $gF$ induces a bijection $G(q)/B(q)\to\mathcal{B}_{r}$ between the two $G$-sets. Thus, by Lemma 3.1(1), the $G$-orbits in ${\mathcal{B}}_{r}\times{\mathcal{B}}_{r}$ correspond bijectively to the set of double coset $B\setminus G/B$. Hence, the the Bruhat decomposition, every $G$-orbit $G.(\mathsf{F},\mathsf{F}^{\prime})$ in ${\mathcal{B}}_{r}\times{\mathcal{B}}_{r}$ has the from
@@
In general, for $\lambda\in\Lambda(n+1,r)$, where $2n+1\leq 2r$, let $P_{\widehat{\lambda}}(q)$ be the standard parabolic subalgebra of ${\text{\rm GL}}_{2r}(\mathbb{F}_{q})$ associated with $\widehat{\lambda}$, consisting of upper quasi-triangular matrices with blocks of sizes $\widehat{\lambda}_{i}$ on the diagonal. Let
@@ 3.4. Parametrizing ${\mathrm{O}}_{2r}(q)$-orbits
${\bf dim}:{\mathcal{F}}_{n,r}^{\jmath}\longrightarrow\widehat{\Lambda}(n+1,r),\;\;F\longmapsto{\bf dim}(F).$ (3.2.1)
Here ${\bf dim}(F):=(\text{\rm dim\,}F_{1}/F_{0},\text{\rm dim\,}F_{2}/F_{1},\ldots,\text{\rm dim\,}F_{n}/F_{n-1},\text{\rm dim\,}F_{n+1}/F_{n},\ldots,\text{\rm dim\,}F_{N}/F_{N-1})$ is called the dimension sequence (or dimension vector), induces a bijective map from ${\mathcal{F}}_{n,r}^{\jmath}/G(q)$ to $\widehat{\Lambda}(n+1,r)$. Thus, the $G$-orbits are indexed by the set $\Lambda(n+1,r)$.
@@
A similar refinement $F_{\!\centerdot\centerdot}^{\prime}$ for $F^{\prime}$ can be defined by swapping the roles of $F$ and $F^{\prime}$.
@@
The dimension vector dim associated with $F_{\!\centerdot\centerdot}$ (see (3.2.1)) defines a matrix
${\mathfrak{m}}(F,F^{\prime})=(a_{i,j})\text{ with }a_{i,j}=\text{\rm dim\,}(F_{i,j}/F_{i,j-1})=\text{\rm dim\,}(F^{\prime}_{j,i}/F^{\prime}_{j,i-1}).$ (3.3.1)
Here the equality follows from the vector space isomorphism $F_{i,j}/F_{i,j-1}\cong\frac{F_{i}\cap F_{j}^{{}^{\prime}}}{F_{i-1}\cap F_{j}^{{}^{\prime}}+F_{i}\cap F_{j-1}^{{}^{\prime}}}\cong F^{\prime}_{j,i}/F^{\prime}_{j,i-1}$ by the Zassenhaus Lemma.
@@
(2) Associated with $A\in\Xi_{N,2r}$, there is a permutation, a distinguished double coset representative, sending $k$ to $i_{k}$, where $(i_{1},i_{2},\cdots,i_{2r})$ is the a permutation of $1,2,\ldots,2r$ obtained by concatenating $I_{1,1}$ with $I_{2,1}$, down column 1 and then down column 2, and so on. See Lemma 2.2 and [8, Exer.8.2(1)]. Then, by definition, $F^{A}_{i}=\langle v_{i_{1}},\cdots,v_{i_{\tilde{\mu}_{i}}}\rangle$, where $\mu=(\mu_{1},\ldots,\mu_{N})=\text{co}(A)$. By part (1), $F^{A}$ is isotropic as it can be obtained by dropping some steps from a corresponding complete isotropic flag.
@@
$A=\begin{pmatrix}1&2&3\\ 3&0&{\bf 3}\\ {\bf 3}&{\bf 2}&{\bf 1}\end{pmatrix}\quad$
@@
Breaking the standard basis $(e_{1},e_{2}\ldots,e_{18})$ into subsequences of lengths $a_{1,1},a_{1,2},a_{1,3},a_{2,1},\ldots,a_{3,3}$ and form the matrix of the subsequences
@@
For vector spaces $U,V$, $n$-step isotropic flags $F,F^{\prime},\in\mathcal{F}_{n,r}^{\jmath}$, and $h\in[1,n]$, we set
@@
$a^{\prime}_{h+1,j}$ $=a_{h+1,j}-(|\textsc{b}^{\prime}_{h,j}|-|\textsc{b}_{h,j}|)+(|\textsc{b}^{\prime}_{h,j-1}|-|\textsc{b}_{h,j-1}|)$
@@
Let $(\mathcal{F}_{n,r}^{\jmath}\times\mathcal{F}_{n,r}^{\jmath})/G(q)$ denote the set of $G(q)$-orbits in $\mathcal{F}_{n,r}^{\jmath}\times\mathcal{F}_{n,r}^{\jmath}$ Then, we obtain a map

### v2

@@ 2. The ${\boldsymbol{q}}$-Schur algebras of type $D$: an algebraic setting
$S_{{\boldsymbol{q}}}^{\kappa}(n,r)=\text{\rm End}_{\check{H}_{\boldsymbol{q}}}\big{(}\bigoplus_{\lambda\in\Lambda(n+1,r)}x_{\lambda}H_{{\boldsymbol{q}},1}|_{\check{H}_{\boldsymbol{q}}}\big{)}$ (2.1.1)
Let ${\boldsymbol{q}}=\upsilon^{2}$ and ${\mathcal{Z}}:=\mathbb{Z}[\upsilon,\upsilon^{-1}]$. By base change to ${\mathcal{Z}}$, we obtain the Hecke algebras ${\mathcal{H}}_{\upsilon,1},\check{\mathcal{H}}_{\upsilon}$ and the $\upsilon$-Schur algebras ${\mathcal{S}}_{\upsilon,1}^{\textsc{b}}(n,r)$ and ${\mathcal{S}}_{\upsilon}^{\kappa}(n,r)$:
${\mathcal{H}}_{\upsilon,1}:=H_{{\boldsymbol{q}},1}\otimes{\mathcal{Z}},\;\;\check{\mathcal{H}}_{\upsilon}:=\check{H}_{\boldsymbol{q}}\otimes{\mathcal{Z}},\;\;{\mathcal{S}}_{\upsilon,1}^{\textsc{b}}(n,r):=S_{{\boldsymbol{q}},1}^{\textsc{b}}(n,r)\otimes_{{\mathcal{A}}}{\mathcal{Z}},\;\;{\mathcal{S}}_{\upsilon}^{\kappa}(n,r)=S_{{\boldsymbol{q}}}^{\kappa}(n,r)\otimes_{{\mathcal{A}}}{\mathcal{Z}}.$ (2.1.2)
Though we will not consider these ${\mathcal{Z}}$-algebras in the rest of the paper, they play important roles for the study of the Schur–Weyl–Hecke duality mentioned in the introduction.
2.2. Identifying $W,W_{\lambda}$ with certain fixed-point subgroups of ${\mathfrak{S}}_{2r}$
@@
Let ${\mathfrak{S}}_{\mu}$ be the Young subgroup of ${\mathfrak{S}}_{2r}$ associated with $\mu$. This is the stabilizer subgroup of the standard Young tabloid [^2] If we identify $\mu$ with its Young diagram consisting of $\mu_{1}$ boxes in row 1, and $\mu_{2}$ boxes in row 2 and so on, then a $\mu$-tableau is obtained by filling the numbers $1,2,\ldots,2r$ into boxed. A $\mu$-tabloid is a row-equivalent class of $\mu$-tableaux. $R^{\mu}=R^{\mu}_{1}\times\cdots\times R_{N}^{\mu}$, where $R^{\mu}_{i}:=[\widetilde{\mu}_{i-1}+1,\widetilde{\mu}_{i-1}+\mu_{i}]$, for all $i=1,2,\ldots,N$. In other words,
@@
consisting of permutations which stabilize each of the subsets $R^{\mu}_{i}$.
@@
^$\widehat{\ }:\Lambda(n+1,r)\longrightarrow\Lambda(2n+1,2r),\;\;\lambda\longmapsto\widehat{\lambda}.$ (2.1.8)
Let $\widehat{\Lambda}(n+1,r)$ denote its image. We now describe the parabolic subgroup $W_{\lambda}$ defined in (2.0.7) in terms of stabilizers of the subsets $R_{i}^{\lambda}:=R_{i}^{\widehat{\lambda}}$, for $1\leq i\leq{\mathsf{m}}$, where ${\mathsf{m}}={\mathsf{m}}(\lambda)$ denotes the maximal index $i$ satisfying $\lambda_{i}\neq 0$.
@@
    Suppose the double coset ${\mathfrak{S}}_{\widehat{\lambda}}w{\mathfrak{S}}_{\widehat{\mu}}$ defines a $(2n+1)\times(2n+1)$ matrix $A=(a_{i,j})$ over $\mathbb{N}$ whose entries sum to $2r$. Then ${\mathfrak{S}}_{\widehat{\lambda}}w{\mathfrak{S}}_{\widehat{\mu}}$ is stabilized by $\theta$ if and only if $a_{i,j}=a_{N+1-i,N+1-j}$ for all $i,j\in[1,N]$ ($N=2n+1$). In particular, for ${\widehat{d}}=\iota(d)\in{\mathfrak{S}}_{2r}$, $\theta$ stabilizes ${\mathfrak{S}}_{\widehat{\lambda}}{\widehat{d}}{\mathfrak{S}}_{\widehat{\mu}}$ and the double coset $W_{\lambda}dW_{\mu}=({\mathfrak{S}}_{\widehat{\lambda}}{\widehat{d}}{\mathfrak{S}}_{\widehat{\mu}})^{\theta}$.
@@
$\text{\rm ro}(A)=\Big{(}\sum_{j=1}^{N}a_{1,j},\sum_{j=1}^{N}a_{2,j},\ldots,\sum_{j=1}^{N}a_{N,j}\Big{)},\quad\text{\rm co}(A)=\Big{(}\sum_{i=1}^{N}a_{i,1},\sum_{i=1}^{N}a_{i,2},\ldots,\sum_{i=1}^{N}a_{i,N}\Big{)}.$ (2.3.2)
@@
Each triple $(\lambda,d,\mu)\in{\mathcal{D}}(n,r)$ defines a natural basis element in $S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)$. We denote this element by $e_{A}$ if $A=\mathfrak{d}(\lambda,d,\mu)$. By definition,
@@
The set $\{e_{A}\mid A\in\Xi_{2n+1,2r}\}$ forms a basis for $S^{\textsc{b}}_{{\boldsymbol{q}},1}(n,r)$ which we call its natural basis.
2.3. Parabolic subgroups of ${\check{W}}$ as stabilizers of tabloids associated with signed compositions
We now extend Lemma 2.2 above to (standard) parabolic subgroups of ${\check{W}}$; compare [19]. Let $\lambda\in\Lambda(n+1,r)$ have maximal index ${\mathsf{m}}$.
@@ 2.3. Parabolic subgroups of ${\check{W}}$ as stabilizers of tabloids associated with signed compositions
(1) The case (2) above shows that distinct tabloids may have the same stabilizer subgroups. This phenomenon is similar to the type $A$ fact that distinct compositions may define the same Young subgroup of the symmetric group.
(2) Consider the parabolic subgroup $W_{\mu}=\langle s_{1},\ldots,s_{r-1}\rangle$ of $W$, where $\mu={(r,0^{n})}$. This is also a parabolic subgroup of $\check{W}$. That is $W_{\mu}=\check{W}_{\mu}$. Clearly, $\varsigma=s_{r}s_{r-1}s_{r}\in{\mathcal{D}}_{\mu,\mu}$, but $s_{r-1}=s_{r}\varsigma s_{r}$ is not distinguished in the double coset $\check{W}_{\mu}(s_{r}ds_{r})\check{W}_{\mu}=\check{W}_{\mu}$.
2.4. The ${\boldsymbol{q}}$-Schur algebras of type $D$
We are now ready to introduce the $q$-Schur algebra of type $D$ in terms of ${\boldsymbol{q}}$-permutation modules of $\check{H}_{\boldsymbol{q}}$ (see [19, (A.3.3)] and compare [11, Def. 2.5]). Let
@@ 2.4. The ${\boldsymbol{q}}$-Schur algebras of type $D$
$S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$ $:=\text{\rm End}_{\check{H}_{\boldsymbol{q}}}\Big{(}\bigoplus_{\alpha\in\Lambda^{\textsc{d}}(n,r)}{\check{x}}_{\alpha}{\check{H}_{\boldsymbol{q}}}\Big{)}\quad(\text{over }{\mathcal{A}}=\mathbb{Z}[{\boldsymbol{q}}]);$
${\mathcal{S}}_{\upsilon}^{\textsc{d}}(n,r)$ $:=\text{\rm End}_{{\check{\mathcal{H}}}_{\upsilon}}\Big{(}\bigoplus_{\alpha\in\Lambda^{\textsc{d}}(n,r)}{\check{x}}_{\alpha}{\check{\mathcal{H}}}_{\upsilon}\Big{)}\quad(\text{over }{\mathcal{Z}}=\mathbb{Z}[\upsilon,\upsilon^{-1}]).$
@@
We will introduce the natural basis for $S_{\boldsymbol{q}}^{\textsc{d}}(n,r)$ and discuss its geometric intepretation in Section 4.
@@
In [4], a geometric framework is introduced for the $q$-Schur algebra $S^{\jmath}_{q}(n,r)$ of type $B$ in equal parameter. We now modify their construction using $(2n+1)$-step filtrations for the $2r$-dimensional space $\mathbb{F}_{q}^{2r}$ (see [14]) and the action by the orthogonal group ${\text{\rm O}}_{2r}(q)$. This results in a $q$-Schur algebra in unequal parameters. Restricting the action to the subgroup ${\text{\rm SO}}_{2r}(q)$ leads to a geometric setting for the $q$-Schur algebra of type $D$; see Theorems 3.10 and 4.15 in the next two sections.
@@ 3.1. A geometric setting for the endomorphism algebra of a permutation module
The $G$-action is extended diagonally to $X\times X$. Consider three projection maps $\pi_{i,j}:X\times X\times X\to X\times X$ sending $(x_{1},x_{2},x_{3})$ to $(x_{i},x_{j})$ for $(i,j)\in\{(1,2),(2,3),(1,3)\}$. Then $\mathfrak{F}_{G}(X\times X,{\mathcal{R}})$ becomes an associative algebra with multiplication defined by $f*f^{\prime}=(\pi_{1,3})_{!}(\pi_{1,2}^{*}f\cdot\pi_{2,3}^{*}f^{\prime})$. In other words,
@@ 3.2. Finite orthogonal groups and flag varieties
We now look at the case, where $G=O_{2r}(q)$ or $SO_{2r}(q)$ is the finite orthogonal or special orthogonal group, and $X=\mathcal{F}_{n,r}^{\jmath}$ is the set of $n$-step filtrations (or partial flag variety) of isotropic subspaces.
@@
This is a group with a split BN-pair, where $B=\widehat{B}(q)\cap{\text{\rm O}}_{2r}(q)$ with $\widehat{B}(q)$ being the Borel subgroup of ${\text{\rm GL}}_{2r}(q)$ consisting of upper triangular matrices and $N=\{\dot{w}\mid w\in W\}$ (see, e.g., [15, p.80]), where
@@ 3.3. Parabolic subgroups of ${\text{\rm O}}_{2r}(q)$
Consider the natural basis $e_{1},e_{2},\ldots,e_{2r}$ for $\mathbb{F}_{q}^{2r}$ and form the (standard) pure isotropic complete flag $\mathsf{F}^{s}$ with
@@
Then we have $B(q)=\text{Stab}_{G(q)}(\mathsf{F}^{s})$. Moreover, the mapping $G(q)\to\mathcal{B}_{r}$ sending $g$ to $gF$ induces a bijection $G(q)/B(q)\to\mathcal{B}_{r}$ between the two $G$-sets. Thus, by Lemma 3.1(1), the $G$-orbits in ${\mathcal{B}}_{r}\times{\mathcal{B}}_{r}$ correspond bijectively to the set of double coset $B\setminus G/B$. Hence, the Bruhat decomposition, every $G$-orbit $G.(\mathsf{F},\mathsf{F}^{\prime})$ in ${\mathcal{B}}_{r}\times{\mathcal{B}}_{r}$ has the from
@@
In general, for $\lambda\in\Lambda(n+1,r)$, where $2n+1\leq 2r$, let $P_{\widehat{\lambda}}(q)$ be the standard parabolic subalgebra of ${\text{\rm GL}}_{2r}(q)$ associated with $\widehat{\lambda}$, consisting of upper quasi-triangular matrices with blocks of sizes $\widehat{\lambda}_{i}$ on the diagonal. Let
@@ 3.4. Parametrizing ${\mathrm{O}}_{2r}(q)$-orbits
${\bf dim}:{\mathcal{F}}_{n,r}^{\jmath}\longrightarrow\widehat{\Lambda}(n+1,r),\;\;F\longmapsto{\bf dim}(F),\;\;\text{where}$ (3.2.1)
${\bf dim}(F):=($ $\text{\rm dim\,}F_{1}/F_{0},\text{\rm dim\,}F_{2}/F_{1},\ldots,\text{\rm dim\,}F_{n}/F_{n-1},\text{\rm dim\,}F_{n+1}/F_{n},\ldots,\text{\rm dim\,}F_{N}/F_{N-1})$
is called the dimension sequence (or dimension vector), induces a bijective map from ${\mathcal{F}}_{n,r}^{\jmath}/G(q)$ to $\widehat{\Lambda}(n+1,r)$. Thus, the $G$-orbits are indexed by the set $\Lambda(n+1,r)$.
@@
For simplicity, we write $F_{\!\centerdot\centerdot}$ for $(F,F^{\prime})_{\!\centerdot\centerdot}$ and $F^{\prime}_{\!\centerdot\centerdot}$ for $(F^{\prime},F)_{\!\centerdot\centerdot}$.
@@
The dimension vector dim associated with $F_{\!\centerdot\centerdot}$ (see (3.2.1)) gives rise to a dimension matrix
${\mathfrak{m}}(F,F^{\prime}):={\bf dim}(F,F^{\prime})_{\!\centerdot\centerdot}=(a_{i,j}),\quad\text{where }a_{ij}=\text{\rm dim\,}(F_{i,j}/F_{i,j-1}).$ (3.3.1)
Note that dim ($\text{\rm dim\,}(F_{i,j}/F_{i,j-1})=\text{\rm dim\,}(F^{\prime}_{j,i}/F^{\prime}_{j,i-1})$, which follows from the vector space isomorphism
$F_{i,j}/F_{i,j-1}\cong\frac{F_{i}\cap F_{j}^{{}^{\prime}}}{F_{i-1}\cap F_{j}^{{}^{\prime}}+F_{i}\cap F_{j-1}^{{}^{\prime}}}\cong F^{\prime}_{j,i}/F^{\prime}_{j,i-1}$
by the Zassenhaus Lemma.
@@
(2) Associated with $A\in\Xi_{N,2r}$, there is a permutation, a distinguished double coset representative, sending $k$ to $i_{k}$, where $(i_{1},i_{2},\cdots,i_{2r})$ is the permutation of $1,2,\ldots,2r$ obtained by concatenating $I_{1,1}$ with $I_{2,1}$, down column 1 and then down column 2, and so on. See Lemma 2.2 and [8, Exer.8.2(1)]. Then, by definition, $F^{A}_{i}=\langle v_{i_{1}},\cdots,v_{i_{\tilde{\mu}_{i}}}\rangle$, where $\mu=(\mu_{1},\ldots,\mu_{N})=\text{co}(A)$. By part (1), $F^{A}$ is isotropic as it can be obtained by dropping some steps from a corresponding complete isotropic flag.
@@
$A=\begin{pmatrix}1&2&3\\ 3&0&{\bf 3}\\ {\bf 3}&{\bf 2}&{\bf 1}\end{pmatrix}.\quad$
@@
Breaking the natural basis $(e_{1},e_{2}\ldots,e_{18})$ into subsequences of lengths $a_{1,1},a_{1,2},a_{1,3},a_{2,1},\ldots,a_{3,3}$ and form the matrix of the subsequences
@@
For vector spaces $U,V$, $n$-step isotropic flags $F,F^{\prime}\in\mathcal{F}_{n,r}^{\jmath}$, and $h\in[1,n]$, we set
@@
We set $E_{ij}^{\theta}=E_{ij}+E_{N+1-i,N+1-j},$ where $E_{ij}$ is the $N\times N$ matrix whose $(i,j)$-entry is $1$ and all other entries are 0.
@@
$a^{\prime}_{h+1,j}$ $=a_{h+1,j}-(|\textsc{b}^{\prime}_{h,j}|-|\textsc{b}_{h,j}|)+(|\textsc{b}^{\prime}_{h,j-1}|-|\textsc{b}_{h,j-1}|).$
@@
Let $(\mathcal{F}_{n,r}^{\jmath}\times\mathcal{F}_{n,r}^{\jmath})/G(q)$ denote the set of $G(q)$-orbits in $\mathcal{F}_{n,r}^{\jmath}\times\mathcal{F}_{n,r}^{\jmath}$. Then, we obtain a map